the sun of scotts age and his mothers age is 45. in 5 years time, three times scotts age less 9 will be the same as his mothers age. find the present ages of scott and his other.
Answer by 303795(602)

S + M = 45
In 5 years time Scott will be S+5 years old and Mum will be M+5
3*(S+5)-9=M+5
3S+6=M+5
M=3S+1
And from the original equation
M=45-S
so
45-S=3S+1
45-S+S=3S+1+S
45=4S+1
44=4S
so Scott is currently 11 and mum is currently 34